OpenStudy (anonymous):

The perimeter of a rectangle is 38m. The length is 10m more than twice the width. Find the demensions. What is the lenght? ?=m Simplify your answer

OpenStudy (anonymous):

6

OpenStudy (anonymous):

sorry, 16m

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

try setting up the equations: you know that the perimeter of a box is twice the length + twice the width

OpenStudy (anonymous):

so let's use w for width and l for length we have that 2w+2l = 38

OpenStudy (anonymous):

but you are told that the length is 10 more than twice the width...so can you write that in equation form? i'll get you started.... length = 2 times width plus 10 (<=== write that in equation form)

OpenStudy (anonymous):

what is the width I got -L+19

OpenStudy (anonymous):

so you have w = -L + 19?

OpenStudy (anonymous):

try plugging in your length...we know length is equal to 16

OpenStudy (anonymous):

what do you get?

OpenStudy (anonymous):

so... w = -(16) + 19 = ?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

correct test your answer...does everything work out?

OpenStudy (anonymous):

remember our original idea....that 2*length + 2*width = 38, your perimeter